OpenStudy (anonymous):

Point E is located at (2, −3) and point F is located at (−2, −1). Find the y value for the point that is 3 over 4 the distance from point E to point F.

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

i bet we can do this

OpenStudy (misty1212):

especially since they only want the y coordinate, not both!

OpenStudy (misty1212):

from \(-3\) to \(-1\) is \(2\) units up three fourths of 2 is \(\frac{3}{4}\times 2=\frac{3}{2}\) or maybe \(1\tfrac{1}{2}\) right ? r

OpenStudy (misty1212):

add that to \(-3\) i.e. compute \(-3+\frac{3}{2}\) and that will be your answer

OpenStudy (anonymous):

-1.5?